# DRCT: Saving Image Super-Resolution away from Information Bottleneck

### v4

@@ 1 Introduction
To address the issue of spatial information loss due to an increased number of network layers, we introduce the Dense-residual-connected Transformer (DRCT), designed to stabilize the forward-propagation process and prevent information bottlenecks. This is achieved by the proposed Swin-Dense-Residual-Connected Block (SDRCB), which incorporates Swin Transformer Layers and transition layers into each Residual Deep feature extraction Group (RDG). Consequently, this approach enhances the receptive field with fewer parameters and a simplified model architecture, thereby resulting in improved performance. The main contributions of this paper are summarised as follows:
@@ 5.1 Network Architecture
Figure 3: The overall architecture of the proposed DRCT and the structure of RDG and SDRCB.
@@
where $H_{DF}(\cdot)$ is the deep feature extraction module and it contains $K$ Residual Deep feature extraction Group (RDG) and single convolution layer $\text{Conv}(\cdot)$ for feature transition. More specifically, intermediate features $\mathbf{F}_{1},\mathbf{F}_{2},\ldots,\mathbf{F}_{K}$ and the output deep feature $\mathbf{F}_{DF}$ are extracted block by block as
@@ 5.2 Deep Feature Extraction
Residual Deep feature extraction Group. In developing of RDG, we take cues from RRDB-Net [52] and RDN [61], employing a residual-dense block (RDB) as the foundational unit for SISR. The reuse of feature maps emerges as the enhanced receptive field in the RDG’s feed-forward mechanism. To expound further, RDG with several SDRCB enhances the capability to integrate information across different scales, thus allowing for a more comprehensive feature extraction. RDG facilitates the information flow within residual group, capturing the local features and spatial information group by group.
@@ 6 Experiment Results
DRCT (Ours) ×$\times 2$ DF2K 38.72 0.9646 34.96 0.9287 32.75 0.9071 34.54 0.9474 40.41 0.9814

### v5

@@ 1 Introduction
To address the issue of spatial information loss due to an increased number of network layers, we introduce the Dense-residual-connected Transformer (DRCT), designed to stabilize the forward-propagation process and prevent information bottlenecks. This is achieved by the proposed Swin-Dense-Residual-Connected Block (SDRCB), which incorporates Swin Transformer Layers and transition layers into each Residual Dense Group (RDG). Consequently, this approach enhances the receptive field with fewer parameters and a simplified model architecture, thereby resulting in improved performance. The main contributions of this paper are summarised as follows:
@@ 5.1 Network Architecture
Figure 3: The overall architecture of the proposed Dense-residual-connected Transformer (DRCT) and the structure of Residual-Dense Group (RDG). Each RDG contains five consecutive Swin-Dense-Residual-Connected Blocks (SDRCBs). By integrating dense-connection [15] into SwinIR [34], the efficiency can be improved for Saving Image Super-resolution away from Information Bottleneck.
@@
where $H_{DF}(\cdot)$ is the deep feature extraction module and it contains $K$ Residual Dense Group (RDG) and single convolution layer $\text{Conv}(\cdot)$ for feature transition. More specifically, intermediate features $\mathbf{F}_{1},\mathbf{F}_{2},\ldots,\mathbf{F}_{K}$ and the output deep feature $\mathbf{F}_{DF}$ are extracted block by block as
@@ 5.2 Deep Feature Extraction
Residual Dense Group. In developing of RDG, we take cues from RRDB-Net [52] and RDN [61], employing a residual-dense block (RDB) as the foundational unit for SISR. The reuse of feature maps emerges as the enhanced receptive field in the RDG’s feed-forward mechanism. To expound further, RDG with several SDRCB enhances the capability to integrate information across different scales, thus allowing for a more comprehensive feature extraction. RDG facilitates the information flow within residual group, capturing the local features and spatial information group by group.
@@ 6 Experiment Results
DRCT (Ours) ×$\times 2$ DF2K 38.62 0.9628 34.84 0.9272 32.62 0.9051 34.44 0.9464 40.31 0.9804